# Covering Complete Geometric Graphs by Monotone Paths

### v1

@@ 5 Concluding remarks
1. For $k\geq 2$, let $f(k,n)$ denote the maximum number of edges of a geometric graph on $n$ vertices that contains no $k$ pairwise disjoint edges. Early results of Erdős [7], Kupitz [11], and Alon and Erdős [2] show that $f(2,n)=n$ and $f(3,n)=O(n)$. The first polynomial upper bound for $f(k,n)$, for $k>3$, was established by Pach and Törőcsik [16], and was improved by Tóth and Valtr [20]; see also [14, Chap. 14]. The current best result, $f(k,n)=O(k^{2}n)$ is due to Tóth [19]. It was used in the proof of Theorem 1.1.
It is conjectured that $f(k,n)=O(kn)$ [5, Chap. 9.5]. If this conjecture is true, our methods would yield an $O(n\log{n})$ (unfortunately, still superlinear) upper bound on the number of monotone matchings sufficient for covering the edge set of $K_{n}[A]$, analogous to covering this set by noncrossing matchings in [3].
@@
On the other hand, it is possible that every complete geometric graph on $n$ vertices can be partitioned into $n/2+o(n)$ plane subgraphs. Pach, Saghafian, and Schnider [15] proved that a complete convex geometric graph on $n$ vertices cannot be decomposed into fewer than $n-1$ plane star-forests.
@@
4. Perhaps the first step toward solving Problem 1 would be to answer the following question.
Problem 5.1.
Does there exist a constant $c>1/2$ with the property that, for infinitely many values of $n$, there are $n$-element point sets $A$ in the plane such that any covering of the edge set of $K_{n}[A]$ by noncrossing paths uses at least $cn$ paths?
@@
We claim that for any finite point set $A$ in general position in the plane, one can cover the edge set of $K_{n}[A]$ by plane Hamiltonian zig-zag paths. Equivalently, for any pair of points $a,b\in A$, the segment $ab$ appears in at least one such path. To see this, assume that the line through $a,b$ separates $A\setminus\{a,b\}$ into two disjoint parts $A_{1},A_{2}$. The edge $ab$ is an edge of the convex hull of $A_{i}\cup\{a,b\}$, for $i=1,2$. Starting with $ab$, we construct a plane zig-zag path that traverses all points in $A_{1}\cup\{a,b\}$, as follows. At the first step, let $b_{1}\in A_{1}$ be the point such that the angle $\angle abb_{1}$ is as small as possible. We continue this process by selecting $b_{i}$ at the $i$-th step such that the angle $\angle b_{i-2}b_{i-1}b_{i}$ is minimized among all remaining points in $A_{1}$, where $b_{0}=b$. It is not hard to verify that once all points in $A_{1}$ are included, the resulting path is a plane zig-zag path. Similarly, starting with $ba$ we can construct a plane zig-zag path that traverses all points in $A_{2}\cup\{a,b\}$. The union of these two paths is a plane zig-zag Hamiltonian path, as claimed.

### v2

@@ 5 Concluding remarks
1. For $k\geq 2$, let $f(k,n)$ denote the maximum number of edges of a geometric graph on $n$ vertices that contains no $k$ pairwise disjoint edges. Early results of Erdős [7], Kupitz [11], and Alon and Erdős [2] show that $f(2,n)=n$ and $f(3,n)=O(n)$. The first polynomial upper bound for $f(k,n)$, for $k>3$, was established by Pach and Törőcsik [15], and was improved by Tóth and Valtr [19]; see also [13, Chap. 14]. The current best result, $f(k,n)=O(k^{2}n)$ is due to Tóth [18].
It is conjectured that $f(k,n)=O(kn)$ [5, Chap. 9.5]. If this conjecture is true, our methods would yield an $O(n\log{n})$ (unfortunately, still superlinear) upper bound on the number of monotone matchings sufficient for covering the edge set of $K_{n}[A]$, analogous to covering this set by crossing-free matchings in [3].
@@
On the other hand, it is possible that every complete geometric graph on $n$ vertices can be partitioned into $n/2+o(n)$ plane subgraphs. Pach, Saghafian, and Schnider [14] proved that a complete convex geometric graph on $n$ vertices cannot be decomposed into fewer than n1n-1 plane star-forests.
Perhaps the first step towards solving Problem 1 would be to answer the following question.
Problem 5.1.
Does there exist a constant $c>1/2$ with the property that, for infinitely many values of $n$, there are $n$-element point sets $A$ in the plane such that every covering of the edge set of $K_{n}[A]$ by crossing-free paths requires at least $cn$ paths?
@@
We show that for any finite point set $A$ in general position in the plane, one can cover the edge set of $K_{n}[A]$ by plane Hamiltonian zig-zag paths. Equivalently, for any pair of points $a,b\in A$, the segment $ab$ appears in at least one such path. To see this, assume that the line through $a,b$ separates $A\setminus\{a,b\}$ into two disjoint parts $A_{1},A_{2}$. The edge $ab$ is an edge of the convex hull of $A_{i}\cup\{a,b\}$, for $i=1,2$. Starting with $ab$, we construct a plane zig-zag path that traverses all points in $A_{1}\cup\{a,b\}$, as follows. At the first step, let $b_{1}\in A_{1}$ be the point such that the angle $\angle abb_{1}$ is as small as possible. We continue this process by selecting $b_{i}$ at the $i$-th step such that the angle $\angle b_{i-2}b_{i-1}b_{i}$ is minimized among all remaining points in $A_{1}$, where $b_{0}=b$. It is not hard to verify that once all points in $A_{1}$ are included, the resulting path is a plane zig-zag path. Similarly, starting with $ba$ we can construct a plane zig-zag path that traverses all points in $A_{2}\cup\{a,b\}$. The union of these two paths is a plane zig-zag Hamiltonian path, as claimed.